OpenStudy (anonymous):

Find all solutions of the equation in the interval [0,2pi). 2sin^2x+3sinx+1=0

OpenStudy (mertsj):

\[(2\sin x+1)(\sin x+1)=0\]

OpenStudy (anonymous):

maybe i dunno

OpenStudy (anonymous):

nevermind completely.

OpenStudy (anonymous):

read it way wrong

OpenStudy (mertsj):

\[\sin x=\frac{-1}{2} sinx=-1\]

OpenStudy (mertsj):

\[x=\frac{7\pi}{6}, \frac{11\pi}{6}, \frac{3\pi}{2}\]